What mass, in grams, of argon gas occupy 125 cm3 at 32.0 oC and 103.5 kPa?


1
Expert's answer
2021-02-26T06:04:53-0500

Given:-

"T=32^oC\Rightarrow32+273\Rightarrow305K"

"V=125cm^3\Rightarrow 0.125L"

"P=103.5kPa\Rightarrow1.0214 atm"

Since, from ideal gas equation

"\Rightarrow" "PV=nRT"

"\Rightarrow" "PV=\dfrac{given mass}{molarmass} \times{R}\times{T}"


"\Rightarrow" "given mass=\dfrac{{P}\times{V}\times{molarmass}}{{R}\times{T}}"

"\Rightarrow" "givenmass=\dfrac{{1.0124}\times{0.125}\times{39.948}}{{0.0821}\times{305}}"


"\Rightarrow givenmass=" "0.2018" "gm"


So, the mass of Argon gas is "0.2018 gm"